OpenStudy (anonymous):

what is the vertex form of y=-x^2-4x+3?

OpenStudy (anonymous):

Hint \[ y=-x^2-4x+3=-x^2-4x+4-4+3? \]

OpenStudy (anonymous):

Further hint \[ y=-x^2-4x+3=-(x^2+4x+4) +4+3 \]

OpenStudy (anonymous):

Remember \[ (a+b)^2 = a^2 + 2 a b + b^2 \]

OpenStudy (anonymous):

Are you still here?